OpenStudy (slickster256):

how would i Determine the answer to (−5) + 4

OpenStudy (slickster256):

@pooja195

OpenStudy (lynfran):

-5+4 it can be written as 4-5 aslo

OpenStudy (♪chibiterasu):

-5 + 4 is the same as 4 - 5

OpenStudy (slickster256):

oh ok thx i kinda forgot over the summer

OpenStudy (♪chibiterasu):

When adding a positive number to a negative number, it goes backwards. 2 + 7 = 9 -2 + 7 = 5

OpenStudy (anonymous):

yup gotta love those negatives